OpenStudy (amenah8):

How would you solve 3x-2y=12 using slope intercept form.

pooja195 (pooja195):

we need to isolate y right? so we need to subtract 3x from both sides now we have \[\huge~\rm-2y=-3x+12\] any cles on what the next step is?

OpenStudy (amenah8):

the next step is to divide the left side by -2.

pooja195 (pooja195):

yes can you please do that? :)

OpenStudy (amenah8):

I got -2y/-2 = 12-3x/-2 and then y = -6 -3x/-2 and the final answer is y = -6 - 3x/2. Is this correct?

pooja195 (pooja195):

-6? where did that come from?

OpenStudy (amenah8):

I got -6 by doing 12/-2.

pooja195 (pooja195):

\[\huge~\rm~\frac{ −2y=−3x+12 }{ -2y=-2y-2y}\]

OpenStudy (amenah8):

I did that except I didn't have it divided by -2y, I had -2 without the y.

OpenStudy (texaschic101):

keep in mind, when dividing by 2 negatives, the result is positive

pooja195 (pooja195):

y=-3/2x-6

OpenStudy (amenah8):

Ah, so that means that the answer would be a positive -6+3x/2.

pooja195 (pooja195):

so yes :) good job! :)

OpenStudy (texaschic101):

y = 3/2x - 6

OpenStudy (texaschic101):

it need to be in y = mx + b form

pooja195 (pooja195):

\(\color{blue}{\text{Originally Posted by}}\) @pooja195 y=-3/2x-6 \(\color{blue}{\text{End of Quote}}\)

OpenStudy (amenah8):

Alright, thanks you two!

OpenStudy (texaschic101):

it is not -3/2....it is 3/2

OpenStudy (amenah8):

Right.